# Modified TSception for Analyzing Driver Drowsiness and Mental Workload from EEG

### v1

@@ Abstract
Driver drowsiness remains a primary cause of traffic accidents, necessitating the development of real-time, reliable detection systems to ensure road safety. This study presents a Modified TSception architecture designed for the robust assessment of driver fatigue using Electroencephalography (EEG). The model introduces a novel hierarchical architecture that surpasses the original TSception by implementing a five-layer temporal refinement strategy to capture multi-scale brain dynamics. A key innovation is the use of Adaptive Average Pooling, which provides the structural flexibility to handle varying EEG input dimensions, and a two-stage fusion mechanism that optimizes the integration of spatiotemporal features for improved stability. When evaluated on the SEED-VIG dataset and compared against established methods—including SVM, Transformer, EEGNet, ConvNeXt, LMDA-Net, and the original TSception—the Modified TSception achieves a comparable accuracy of 83.46% (vs. 83.15% for the original). Critically, the proposed model exhibits a substantially reduced confidence interval (0.24 vs. 0.36), signifying a marked improvement in performance stability. Furthermore, the architecture’s generalizability is validated on the STEW mental workload dataset, where it achieves state-of-the-art results with 95.93% and 95.35% accuracy for 2-class and 3-class classification, respectively. These improvements in consistency and cross-task generalizability underscore the effectiveness of the proposed modifications for reliable EEG-based monitoring of drowsiness and mental workload.
@@ 2 Background
Recent advancements in driver drowsiness detection have shifted from traditional ML (SVM, KNN) to DL models like EEGNet and the original TSception. While these models excel at capturing complex patterns, they often exhibit high performance volatility. For instance, a model may achieve 90% accuracy on one subject but drop to 70% on another, leading to wide CIs that make them unreliable for real-world safety deployment. Furthermore, many current methodologies are ’task-specific,’ failing to generalize when the cognitive demand shifts from fatigue to high mental workload.
@@ 3.3 Layer Architecture of Modified TSception
The inception window is $[0.5,0.25,0.125]$. The extracted feature is activated by LeakyReLU, and down-sampled using Average Pooling (AP), except for layer 3 which uses Adaptive Average Pooling (ADP). The output is concatenated and batch normalized (${BN}_{t}$). The temporal layer output is:
@@ 3.4 Comparison of Architectural Enhancements
While both TSception and Modified TSception leverage parallel spatiotemporal pathways, proposed model introduces significant refinements. By expanding to five Tception layers and replacing fixed pooling with ADP, the model achieves better scale invariance. The introduction of the $1\times 1$ pointwise fusion layer and a deeper FC network (two hidden layers with dropout) further enhances the model’s ability to extract stable, high-level features, directly improving generalizability across diverse cognitive tasks.
@@ 4.1 Experimental Data
SEED-VIG: Driver Drowsiness Dataset: The SEED-VIG dataset (zheng2017multimodal) is an open-source resource for investigating driver vigilance. It comprises EEG recordings from 23 participants (12 males and 11 females) with a mean age of approximately 23.3 years (SD: 1.4 years) engaged in a simulated driving task designed to mimic real-world conditions. Data were acquired using a 17-channel montage focusing on temporal (FT7, FT8, T7, T8, TP7, TP8) and posterior (CP1, CP2, P1, PZ, P2, PO3, POZ, PO4, O1, OZ, O2) regions at a 1000 Hz resolution. Drowsiness was quantified via the PERCLOS (percentage of eyelid closure) metric, with a threshold of 0.5 defining “awake” versus “drowsy” states. To isolate cognitive features and remove physiological noise, signals were band-pass filtered between 1 and 75 Hz using a zero-phase Butterworth filter. Data were down-sampled from 1000 to 200 Hz to streamline computation while maintaining sufficient resolution for temporal convolution. Continuous data were segmented into one-second non-overlapping epochs, resulting in a standardized input tensor of (1, 17, 200). The resulting 40,710 samples were split into training (70%), validation (15%), and test (15%) sets.
STEW: Mental Workload Dataset: To validate cross-task efficacy, the STEW dataset (lim2018stew) was utilized, which focuses on multitasking workload. Raw EEG data from 48 subjects (aged 18–40) were recorded during the simultaneous capacity (SIMKAP) test. Recordings were performed using an Emotiv EPOC headset (14 channels, 128 Hz sampling rate). The protocol included a 2.5-minute resting state (“No task”) and a 2.5-minute workload condition. Preprocessing for the STEW dataset involved several stages to mitigate the higher noise levels associated with consumer-grade headsets. After initial band-pass filtering, Independent Component Analysis (ICA) was performed to decompose the signal into statistically independent components, allowing for the identification and rejection of components corresponding to eye blinks, lateral eye movements, and temporal muscle activity. Furthermore, epochs were manually inspected and removed as ‘bad epochs’ before the data were min-max scaled and segmented with a 0.5-second overlap to enhance the training set size. The final data tensor dimensions were (1, 14, 128) across 26,910 samples.
@@ 2 Class 2 Class 3 Class
Table 1 and the corresponding bar chart (Fig. 2) present a comparative performance analysis across the SEED-VIG and STEW datasets. Accuracy is reported alongside 95% CIs to ensure a robust comparison. In a practical setting, a narrower CI implies that the system is more likely to maintain its advertised performance level regardless of the specific user’s brainwave patterns, thereby reducing the risk of ‘false negatives’ in subjects who deviate from the population average.

### v2

@@ Abstract
Driver drowsiness is a leading cause of traffic accidents, necessitating real-time, reliable detection systems to ensure road safety. This study proposes a Modified TSception architecture for robust assessment of driver fatigue and mental workload using Electroencephalography (EEG). The model introduces a five-layer hierarchical temporal refinement strategy to capture multi-scale brain dynamics, surpassing the original TSception’s three-layer approach. Key innovations include the use of Adaptive Average Pooling (ADP) for structural flexibility across varying EEG dimensions and a two-stage fusion mechanism to optimize spatiotemporal feature integration for improved stability. Evaluated on the SEED-VIG dataset, the Modified TSception achieves 83.46% accuracy, comparable to the original model (83.15%), but with a significantly reduced confidence interval (0.24 vs. 0.36), indicating better performance stability. The architecture’s generalizability was further validated on the STEW mental workload dataset, achieving state-of-the-art accuracies of 95.93% and 95.35% for 2-class and 3-class classification, respectively. These results show that the proposed modifications improve consistency and cross-task generalizability, making the model a reliable framework for EEG-based safety monitoring.
@@ 2 Background
Recent advancements in driver drowsiness detection have shifted from traditional ML (SVM, KNN) to DL models like EEGNet and the original TSception. While these models excel at capturing complex patterns, they often exhibit high performance volatility. For instance, a model may achieve 90% accuracy on one subject but drop to 70% on another, leading to wide CIs that make them unreliable for real-world safety deployment. Furthermore, many current methodologies are ’task-specific’, failing to generalize when the cognitive demand shifts from fatigue to high mental workload.
@@ 3.3 Layer Architecture of Modified TSception
The inception window is $[0.5,0.25,0.125]$. The extracted feature is activated by LeakyReLU, and down-sampled using AP, except for layer 3 which uses ADP. The output is concatenated and batch normalized (${BN}_{t}$). The temporal layer output is:
@@ 3.4 Comparison of Architectural Enhancements
While both TSception and Modified TSception leverage parallel spatiotemporal pathways, the proposed model introduces significant refinements. By expanding to five Tception layers and replacing fixed pooling with ADP, the model achieves better scale invariance. The introduction of the $1\times 1$ pointwise fusion layer and a deeper FC network (two hidden layers with dropout) further enhances the model’s ability to extract stable, high-level features, directly improving generalizability across diverse cognitive tasks. Compared to established architectures like EEGNet and the original TSception, the Modified TSception offers three distinct advantages:
Scale Invariance: While EEGNet utilizes fixed-size kernels that may fail to capture transient bursts across different sampling rates, our five-layer hierarchy allows for simultaneous analysis of both sustained rhythmic activity and rapid, non-stationary EEG spindles.
Hardware Agnosticism: The transition from fixed AP to ADP removes the dependency on specific input dimensions. This allows the model to be deployed across different EEG headsets (e.g., research-grade vs. consumer-grade) without redesigning the network architecture.
Enhanced Stability: The two-stage fusion mechanism specifically addresses the high inter-subject variability found in traditional CNNs. By refining spatial features with pointwise convolutions before final classification, the model achieves a more stable feature representation, as evidenced by the significant reduction in performance variance across different subjects.
@@ 4.1 Experimental Data
SEED-VIG (Driver Drowsiness Dataset): The SEED-VIG dataset (zheng2017multimodal) is an open-source resource for investigating driver vigilance. It comprises EEG recordings from 23 participants (12 males and 11 females) with a mean age of approximately 23.3 years (SD: 1.4 years) engaged in a simulated driving task designed to mimic real-world conditions. Data were acquired using a 17-channel montage focusing on temporal (FT7, FT8, T7, T8, TP7, TP8) and posterior (CP1, CP2, P1, PZ, P2, PO3, POZ, PO4, O1, OZ, O2) regions at a 1000 Hz resolution. Drowsiness was quantified via the PERCLOS (percentage of eyelid closure) metric, with a threshold of 0.5 defining “awake” versus “drowsy” states. To isolate cognitive features and remove physiological noise, signals were band-pass filtered between 1 and 75 Hz using a zero-phase Butterworth filter. Data were down-sampled from 1000 to 200 Hz to streamline computation while maintaining sufficient resolution for temporal convolution. Continuous data were segmented into one-second non-overlapping epochs, resulting in a standardized input tensor of (1, 17, 200). The resulting 40,710 samples were split into training (70%), validation (15%), and test (15%) sets.
STEW (Mental Workload Dataset): To validate cross-task efficacy, the STEW dataset (lim2018stew) was utilized, which focuses on multitasking workload. Raw EEG data from 48 subjects (aged 18–40) were recorded during the simultaneous capacity (SIMKAP) test. Recordings were performed using an Emotiv EPOC headset (14 channels, 128 Hz sampling rate). The protocol included a 2.5-minute resting state (“No task”) and a 2.5-minute workload condition. Preprocessing for the STEW dataset involved several stages to mitigate the higher noise levels associated with consumer-grade headsets. After initial band-pass filtering, Independent Component Analysis (ICA) was performed to decompose the signal into statistically independent components, allowing for the identification and rejection of components corresponding to eye blinks, lateral eye movements, and temporal muscle activity. Furthermore, epochs were manually inspected and removed as ‘bad epochs’ before the data were min-max scaled and segmented with a 0.5-second overlap to enhance the training set size. The final data tensor dimensions were (1, 14, 128) across 26,910 samples.
@@ 2 Class 2 Class 3 Class
Table 1 and the corresponding bar chart (Fig. 2) present a comparative performance analysis across the SEED-VIG and STEW datasets. Accuracy is reported alongside 95% CIs to ensure a robust comparison. In a practical setting, a narrower CI implies that the system is more likely to maintain its advertised performance level regardless of the specific user’s brainwave patterns, thereby reducing the risk of ’false negatives’ in subjects who deviate from the population average.